OpenStudy (anonymous):

Can you show in steps how to complete this question ? 1 1/2= 4h+ 2/3

OpenStudy (anonymous):

assuming the original question looks like this: \[1 \frac{1}{2} = 4h + \frac{2}{3} \] change all fractions to improper fractions, then rearrange the equation to isolate the 'h' term: \[4h = \frac{3}{2} - \frac{2}{3}\] \[h = \frac{3^2 - 2^2}{6} \times \frac{1}{4}\] you should be able to take it from here.